# On the Moduli Space of Discrete Morse Functions

### v1

@@ 3. The moduli space of discrete Morse functions
One central point of this work is the interplay between abstract cell complexes and geometric cell complexes. Although discrete Morse theory excels on abstract cell complexes, we need to include geometric CW complexes for the comparison to smooth Morse theory in Section 4. Hence, on one hand we need to treat discrete Morse functions on geometric CW complexes and on the other hand we want to treat discrete Morse functions on abstract simplicial complex. We consider both points of view on regular CW complexes as a middle ground, since homotopy types of regular CW complexes uniquely determined by their face posets i.e. their associated abstract CW complexes. Nonetheless, it is important to point out that even though the face poset determines its corresponding regular CW complex up to homotopy, geometric cellular morphisms of regular CW complexes are more general than morphisms of the corresponding abstract CW complexes.
The definition of discrete Morse functions gives rise to a subdivision of the space of discrete functions induced by an oriented hyperplane arrangement in $\mathbb{R}^{X}$:
We observe that each point $p\in\mathbb{R}^{X}$ represents a discrete function $f\colon X\rightarrow\mathbb{R}$ by $f(\sigma)=p_{\sigma}$. The conditions (1) and (2) (and (3)) of Definition 3.2 impose conditions on the components of a discrete function $p\in\mathbb{R}^{X}$ whenever there is a directed edge between the corresponding cells in the Hasse diagram. That is, for the question whether $p$ represents a discrete Morse function it is relevant whether the inequality $p_{\sigma}<p_{\tau}$ holds whenever $\sigma\subset\tau$ is represented by a directed edge in $D(X)$. This information is equivalent to the question on which side of the hyperplane $H_{\sigma}^{\tau}$ in $\mathbb{R}^{X}$ given by $x_{\sigma}=x_{\tau}$ the point $p$ lies. Hence, the conditions (1) and (2) (and (3)) of Definition 3.2 can be checked by analyzing the position of $p$ relative to the hyperplane arrangement $\mathcal{A}$ given by the equations $x_{\sigma}=x_{\tau}$ for all covering face relations $\sigma\subset\tau$ in $X$. We define the following oriented hyperplane arrangement:
Definition 3.9.
Let $E$ be the set of codimension 1 face relations in $X$. We define the hyperplane arrangement $\mathcal{A}(X)\coloneqq\{H_{\sigma}^{\tau}|\sigma\subset\tau\in E\}$ where $H_{\sigma}^{\tau}$ is given by the equation $x_{\sigma}=x_{\tau}$. We observe that the assignment $(\sigma\subset\tau)\mapsto H_{\sigma}^{\tau}$ defines a bijection $E\cong\mathcal{A}$. Then let $\mathcal{L}(X)\coloneqq\{+,-,0\}^{E}$ be the sign vectors for the oriented hyperplane arrangement $\mathcal{A}^{+}(X)$ obtained by orienting $\mathcal{A}(X)$ in the following way: For each hyperplane $H_{\sigma}^{\tau}$ we define the open half-space defined by $x_{\sigma}<x_{\tau}$ to be the positive side of $H_{\sigma}^{\tau}$, the open half-space given by $x_{\sigma}>x_{\tau}$ to be the negative side of $H_{\sigma}^{\tau}$, and on $H_{\sigma}^{\tau}$ itself we have the value 0. This way, $\mathcal{L}(X)$ is the set of all covectors of the oriented hyperplane arrangement $\mathcal{A}^{+}(X)$. We call $\mathcal{A}^{+}(X)$ the Morse arrangement on $\mathbb{R}^{X}$, respectively on $X$.
For any covector $x\in\mathcal{L}(X)$ we define the sets $x_{+}\coloneqq\{e\in E|x_{e}=+\}$, $x_{-}\coloneqq\{e\in E|x_{e}=-\}$, and $x_{0}\coloneqq\{e\in E|x_{e}=0\}$.
Remark 3.10.
@@
The index set $E$ of $\mathcal{L}(X)$ is by construction the set of edges of $D(X)$. Hence, we will use the two different points of view interchangeably without explicitly mentioning it in the notation from now on. Moreover, we drop the $+$ in the notation since there is only one orientation we are interested in. Furthermore, we will denote $\mathcal{A}(X)$ and $\mathcal{L}(X)$ just by $\mathcal{A},\mathcal{L}$, respectively, when there are not multiple complexes to be confused with.
@@
We interpret an element $x\in\mathcal{L}$ as an equivalence class of discrete functions on some regular CW complex $X$. For two functions $p,q\in\mathbb{R}^{X}$ we define $p\sim q$ if and only if for all hyperplanes $H$, the function $p$ lies on the positive side of $H$ if and only if $q$ lies on the positive side of $H$. It turns out that this way the regions which contain discrete Morse functions are identified with their equivalence classes given by inducing the same acyclic matching. The positive region, i.e. the region on the positive side of all hyperplanes, turns out to be exactly the set of all critical discrete Morse functions on $X$. A point $p\in\mathbb{R}^{X}$ being inside a hyperplane $H_{\sigma}^{\tau}$ or on the negative side of $H_{\sigma}^{\tau}$ indicates that if $p$ represents a discrete Morse function, then the cells $\sigma$ and $\tau$ are matched by $p$. Moreover, a function $p\in\mathbb{R}$ is a discrete Morse function if and only if all functions in the same region as $p$ are.
@@
We call any region of $\mathcal{A}$ that consists of discrete Morse functions a Morse region. We call the union of all Morse regions of $\mathcal{A}$ the space of discrete Morse functions on $X$ and denote it by $\mathcal{M}(X)$. We call the unique region $R_{cr}(X)$ that only contains critical discrete Morse functions, i.e. discrete Morse functions for which every cell is critical, the critical region of $\mathcal{A}$. Moreover, we define the essential space of discrete Morse functions $\mathcal{M}_{ess}(X)$ as the intersection $\mathcal{M}_{ess}(X)\coloneqq\overline{R_{cr}(X)}\cap\mathcal{M}(X)$. We define the space of Morse–Benedetti functions $\mathcal{MB}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy monotonicity, semi-injectivity and generacy. We define the space of weak Morse–Benedetti funtcions $\mathcal{MB}_{w}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy semi-injectivity and generacy.
@@
The reflection arrangement associated to $\mathcal{A}(X)$ for a regular CW complex $X$ is the braid arrangement associated to the symmetric group $\mathfrak{S}_{\lvert X\rvert}$. For arbitrary CW complexes, the story is less well behaved because one has a priori no combinatorial control over cover relations being regular.
@@
Figure 1. The Morse arrangement in the space of discrete functions on $I$
@@
Let $X$ be a regular CW complex. Then we have $\mathcal{MB}(X)\subset\overline{R_{cr}(X)}$. Moreover, $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$ and $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{ess}(X)$.
@@
Due to generacy, weak Morse–Benedetti functions cannot have the same value on different cells unless such cells are faces of one another. By semi-injectivity, this can only happen if the two cells are faces of codimension one, i.e. if they are matched. Hence, discrete Morse functions can only fail to be weak Morse–Benedetti functions on intersections of at least one hyperplane of $\mathcal{H}\setminus\mathcal{A}$. Thus, $\mathcal{M}(X)\setminus\mathcal{MB}_{w}(X)$ is contained in the first stratum of the stratification induced by $\mathcal{H}(X)$ which means that $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$. It follows that $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{ess}(X)$ by restricting the stratification induced by $\mathcal{H}$ to $\mathcal{M}_{ess}(X)$. ∎
Having established that the space of discrete Morse functions is homotopically trivial, we are going to investigate its combinatorial and geometrical properties. Since properties of objects are ideally investigated via morphisms, we take a look at meaningful notions of morphisms in this context.
@@
Let $X,Y$ be regular CW complexes. We say that a set of 0-cells $\{\sigma_{i}\}\subset X$ spans a cell $\tau\in X$ if $\{\sigma_{i}\}=(\partial\bar{\tau})_{0}$.
A continuous map $\varphi\colon X\rightarrow Y$ is called cellular if it maps the $n$-skeleton $X_{n}$ to the $n$-skeleton $Y_{n}$. A map $\varphi\colon X\rightarrow Y$ is called order-preserving if it corresponds to an order-preserving map $D(\varphi)\colon D(X)\rightarrow D(Y)$. A map $\varphi\colon X\rightarrow Y$ is called simplicial if
  • it maps 0-cells to 0-cells and
  • for any set of 0-cells $\{\sigma_{i}\}\subset X$ such that $\{\sigma_{i}\}\subset X$ spans an $n$-cell for some $n$, the set $\{\varphi(\sigma_{i})\}\subset Y$ spans a $k$-cell for some $k\leq n$.
@@
We remark that in Definition 3.16 cellular maps belong the the framework of of geometric CW complexes, whereas order-preserving maps and simplicial maps belong to the realm of abstract CW complexes.
There is a more general notion of equivalence of CW complexes in the literature:
Definition 3.18 ([BH04, Definition 6.32]).
@@
The complexes $X$ and $X^{\prime}$ are called cell equivalent if and only if there is a homotopy equivalence $h\colon X\rightarrow X^{\prime}$ with the property that there is a bijective correspondence between cells in $X$ and cells in $X^{\prime}$ such that if $e\subset X$ corresponds to $e^{\prime}\subset X^{\prime}$, then $h$ maps $X(e)$ to $X^{\prime}(e^{\prime})$ and is a homotopy equivalence of these subcomplexes.
@@
Let $X$ and $Y$ be CW complexes. A map $\varphi\colon X\rightarrow Y$ is called non-degenerate if it maps 0-cells to 0-cells, and induces a map $D(\varphi)\colon D(X)\rightarrow D(Y)$ that preserves cover relations, and $\varphi$ induces homotopy equivalences $X(\sigma)\simeq Y(\varphi(\sigma))$ for all $\sigma\in X$.
Remark 3.20.
It is straightforward that for regular CW complexes, cell equivalences induce simplicial equivalences. Moreover, a map $\varphi$ between regular CW complexes is non-degenerate if and only if it maps 0-cells to 0-cells and $D(\varphi)$ preserves cover relations, i.e. face relations of codimension one. In more generality, injective non-degenerate maps between CW complexes are cell equivalences onto their image.
For the context of discrete Morse theory, we prefer the point of view of simplicial maps between regular CW complexes because they focus more on the combinatorial structure at hand. Nonetheless, in the context of arbitrary CW complexes one has to work with cell equivalences because they keep track of attaching maps being irregular.
Furthermore, cell equivalences induce isomorphisms of the corresponding face posets. This way, cell equivalences and non-degenerate maps are the connecting piece between geometric and abstract morphisms of regular CW complexes.
Proposition 3.21.
A map $\varphi\colon X\rightarrow Y$ between regular CW complexes is simplicial if and only if it is cellular and order preserving.
@@
  1. $\Rightarrow$
    Let $\sigma\subseteq\tau$ be a face relation in $D(X)$. The vertices $\sigma_{0}$ form a subset of the vertices $\tau_{0}$. This implies $\varphi(\sigma_{0})\subseteq\varphi(\tau_{0})$. Since $\varphi$ is simplicial, it follows that $\varphi(\sigma_{0})$ spans a cell of dimension less or equal to $\operatorname{dim}\tau$. Thus, $\varphi$ is order-preserving. Moreover, the $\operatorname{dim}\sigma$-skeleton of $X$ is mapped to the $\operatorname{dim}\sigma$-skeleton of $Y$.
  2. $\Leftarrow$
    The property of $\varphi$ being cellular in particular implies that $\varphi$ preserves the 0-skeleton, i.e. maps 0-cells to 0-cells. It follows inductively that for a set of 0-cells $\{\sigma_{i}\}\subset X$, the set $\{\varphi(\sigma_{i})\}\subset Y$ spans a $k$-cell if $\{\sigma_{i}\}\subset X$ spans a $n$-cell for some $k\leq n$. Hence, $\varphi$ is simplicial.
Proposition 3.22.
Non-degenerate maps of regular CW complexes are simplicial and map $n$-cells to $n$-cells. Simplicial maps of CW complexes that map $n$-cells to $n$-cells are non-degenerate.
Proof.
Since in regular CW complexes the combinatorial structure of all cells is given by chains of inclusions of the corresponding faces, the first statement follows by induction over chains of non-trivial inclusions of faces of codimension one.
For the second statement, we observe that simplicial maps in particular map 0-simplices to 0-simplices. For the preservation of cover relations we observe that $n$-cells of regular CW complexes are characterized by chains of cover relations of their faces of length $n+1$. Since these chains need to preserve their length under simplicial maps that map $n$-cells to $n$-cells, cover relations also need to be preserved because otherwise the chains would get shorter. ∎
Lemma 3.23.
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map. Then $\varphi$ induces a map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ that maps $\mathcal{A}^{+}(X)$ to $\mathcal{A}^{+}(Y)$. If $\varphi$ is injective, then $\varphi_{*}$ embeds $\mathcal{A}^{+}(X)$ into $\mathcal{A}^{+}(Y)$.
Proof.
By Proposition 3.22, $\varphi$ maps $n$-cells to $n$-cells and, therefore, induces a map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$. Since $\varphi$ preserves cover relations in $D(X)$ and subcomplexes up to homotopy, $\varphi$ preserves the property of cover relations being regular. Hence, $\varphi_{*}$ maps $\mathcal{A}^{+}(X)$ to $\mathcal{A}^{+}(Y)$. If $\varphi$ is injective, so is $D(\varphi)$ and it follows that $\varphi_{*}$ embeds $\mathcal{A}^{+}(X)$ into $\mathcal{A}^{+}(Y)$. ∎
@@ 2. $\Leftarrow$
In fact, if $\varphi$ is a degenerate simplicial map, $D(\varphi)$ does not preserve cover relations. Since $D(\varphi)$ is a poset map between finite posets, the fiber $D(\varphi)^{-1}(y)$ of a point $y\in D(Y)$ is either empty or an interval in $D(X)$. Intervals in $D(X)$ correspond to subspaces in $\mathbb{R}^{X}$ which means that $\varphi_{*}$ collapses the subspace corresponding to $D(\varphi)^{-1}(y)$ to one coordinate corresponding to $y$. Geometrically, this corresponds to collapsing444in the sense of collapsing a topological space, not necessarily in the sense of simple collapses the corresponding subcomplex $U\subset X$ to a point.
@@
  • $\operatorname{Face}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\subset\sigma\text{ cover relation in }D(X)\}$, and
  • $\operatorname{Coface}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\supset\sigma\text{ cover relation in }D(X)\}$.

### v2

@@ 3. The Space of Discrete Morse Functions
The index set $E$ of $\mathcal{S}(X)$ is by construction the set of edges of $D(X)$. Hence, we will use the two different points of view interchangeably without explicitly mentioning it in the notation from now on. Moreover, we drop the ++ in the notation since there is only one orientation we are interested in. Furthermore, we will denote $\mathcal{A}(X)$ and $\mathcal{S}(X)$ just by $\mathcal{A},\mathcal{S}$, respectively, when there are not multiple complexes to be confused with.
@@
We interpret an element $x\in\mathcal{S}$ as an equivalence class of discrete functions on a given regular CW complex $X$. For two functions $p,q\in\mathbb{R}^{X}$ we define $p\sim q$ if and only if for all hyperplanes $H\in\mathcal{A}(X)$, the function $p$ lies on the positive side of $H$ if and only if $q$ lies on the positive side of $H$. It turns out that this way the regions that contain discrete Morse functions can be identified with their Forman equivalence classes, that is, the ones that induce the same acyclic matching. The positive region, i.e. the region on the positive side of all hyperplanes, turns out to be exactly the set of all critical discrete Morse functions on $X$. A point $p\in\mathbb{R}^{X}$ being inside a hyperplane $H_{\sigma}^{\tau}$ or on the negative side of $H_{\sigma}^{\tau}$ indicates that if $p$ represents a discrete Morse function, then the cells $\sigma$ and $\tau$ are matched by $p$. Moreover, a function $p\in\mathbb{R}^{X}$ is a discrete Morse function if and only if all functions in the same region as $p$ are.
@@
We call any region of $\mathcal{A}$ that consists of discrete Morse functions a Morse region. We call the union of all Morse regions of $\mathcal{A}$ the space of discrete Morse functions on $X$ and denote it by $\mathcal{M}(X)$. We call the unique region $R_{cr}(X)$ that only contains critical discrete Morse functions, i.e. discrete Morse functions for which every cell is critical, the critical region of $\mathcal{A}$. Moreover, we define the space of discrete Morse filtrations [^5] Lying in the closure of the critical region ensures the discrete function to be weakly increasing. This, in turn, ensures that for every cell the function value coincides with its filtration value in the sublevel filtration. $\mathcal{M}_{fil}(X)$ as the intersection $\mathcal{M}_{fil}(X)\coloneqq\overline{R_{cr}(X)}\cap\mathcal{M}(X)$. We define the space of Morse–Benedetti functions $\mathcal{MB}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy monotonicity, semi-injectivity and genericity. We define the space of weak Morse–Benedetti funtcions $\mathcal{MB}_{w}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy semi-injectivity and genericity.
@@
The reflection arrangement associated to $\mathcal{A}(X)$ [^6] I.e.  the smallest reflection arrangement that contains $\mathcal{A}(X)$. for a connected regular CW complex $X$ is the braid arrangement associated to the symmetric group $\mathfrak{S}_{\lvert X\rvert}$. For arbitrary CW complexes, the story is less well behaved because one has a priori no combinatorial control over cover relations being regular.
@@
Figure 3. The Morse arrangement in the space of discrete functions on II
@@
Let $X$ be a regular CW complex. Then we have $\mathcal{MB}(X)\subset\overline{R_{cr}(X)}$. Moreover, $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$ and $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{fil}(X)$.
@@
Due to genericity, weak Morse–Benedetti functions cannot have the same value on different cells unless such cells are faces of one another. By semi-injectivity, this can only happen if the two cells are faces of codimension one, i.e. if they are matched. Hence, discrete Morse functions can only fail to be weak Morse–Benedetti functions on intersections of at least one hyperplane of $\mathcal{H}\setminus\mathcal{A}$. Thus, $\mathcal{M}(X)\setminus\mathcal{MB}_{w}(X)$ is contained in the first stratum of the stratification induced by $\mathcal{H}(X)$ which means that $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$. It follows that $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{fil}(X)$ by restricting the stratification [^8] Any hyperplane arrangement induces a stratification of the ambient vector space where the $k$th-stratum is given by the union of all intersections of $k$ hyperplanes. induced by $\mathcal{H}$ to $\mathcal{M}_{fil}(X)$. ∎
Having established that the space of discrete Morse functions is homotopically trivial, we are going to investigate its combinatorial and geometrical properties. Since properties of objects are ideally investigated via morphisms, we take a look at meaningful notions of morphisms in this context. The goal is to identify morphisms between CW complexes that induce a meaningful notion of morphisms between spaces of discrete Morse functions. Such a notion of morphisms $\varphi\colon X\rightarrow Y$ should
  1. (1)
    induce maps on the spaces of discrete functions,
  2. (2)
    preserve regularity of face relations, and
  3. (3)
    induce maps of hyperplane arrangements that are compatible with the orientation.
We address these points in the order mentioned above, starting with a notion of morphism that induces maps on spaces of discrete functions.
@@
Let $X,Y$ be CW complexes.
A continuous map $\varphi\colon X\rightarrow Y$ is called cellular if it maps the $n$-skeleton $X_{n}$ to the $n$-skeleton $Y_{n}$.
A continuous map $\varphi\colon X\rightarrow Y$ is called order-preserving if it induces an order-preserving map $F(\varphi)\colon F(X)\rightarrow F(Y)$, i.e. if for any cell $\sigma\in X$, there is a cell $\sigma^{\prime}\in Y$, such that $f(\sigma)\subseteq\sigma^{\prime}$ and for cells $\sigma\subset\tau\in X$ we have $f(\sigma)\subset f(\tau)\in Y$.
We call a map $\varphi\colon X\rightarrow Y$ simplicial if it is cellular and order preserving.
@@
One central point of this work is the interplay between CW complexes and their face posets. Although discrete Morse theory excels on simplicial complexes, we need to include CW complexes for the comparison to smooth Morse theory in Section 4.1. Hence, on one hand we need to treat discrete Morse functions on CW complexes and on the other hand we want to treat discrete Morse functions on abstract simplicial complexes. We consider both points of view on regular CW complexes as a middle ground, since homotopy types of regular CW complexes are uniquely determined by their face posets. Nonetheless, it is important to point out that even though the face poset determines its corresponding regular CW complex up to homotopy, geometric cellular morphisms of regular CW complexes are more general than morphisms of the corresponding face posets.
We also remark that in Definition 3.16 cellular maps belong to the framework of CW complexes, whereas order-preserving maps are closely related to face posets. We introduce a notion of simplicial maps as a middle ground between the two.
Example 3.18.
In order to see that cellular maps do not necessarily induce maps on face posets, consider a regular CW decomposition of the circle $S^{1}$ in two 0-cells $a,b$ and two 1-cells $A,B$. Let a continuous map $f\colon S^{1}\rightarrow S^{1}$ be given by $f(a)=f(b)=f(A)=a$ and $f$ is defined on $B$ by winding around the circle twice. Then $f$ does not induce a map on face posets because the image $f(B)$ contains multiple cells, in fact all cells of $S^{1}$. Nonetheless, $f$ maps the 0-skeleton to the 0-skeleton and the 1-skeleton to the 1-skeleton. Since $f$ winds around the circle twice, the map $f$ is also not homotopic to any map that could induce a map on face posets.
Having established notions of morphisms that induce maps on face posets, and thus induce maps of spaces of discrete functions in the sense of partially ordered coordinate spaces, we continue with notions of morphisms that preserve regularity of face relations. In order to do so, we focus on two approaches, which we then combine for regular CW complexes. The first approach uses the combinatorial structure of the faces in the boundary of cells. The second one uses the notion of cell equivalences, which combines homotopy and cell structure.
Definition 3.19.
Let $X$ and $Y$ be CW complexes. We say that a cell $\tau\in X$ is spanned by a set of 0-cells $\{\sigma_{i}\}\subset X$ if $\{\sigma_{i}\}=(\partial\tau)_{0}$. A simplicial map $\varphi\colon X\rightarrow Y$ is called non-degenerate if for any set of 0-cells $\{\sigma_{i}\}\subset X$ such that an $n$-cell $\sigma$ is spanned by $\{\sigma_{i}\}\subset X$ for some $n$, the minimal $k$-cell $\tau$ such that $\varphi(\sigma)\subseteq\tau$ for some $k\leq n$ is spanned by the set $\{\varphi(\sigma_{i})\}\subset Y$.
Example 3.20.
We consider some examples and non-examples for non-degenerate maps:
  1. (1)
    Inclusion of subcomplexes $X\hookrightarrow Y$ are non-degenerate.
  2. (2)
    Let $A\subset X$ be a subcomplex. Then the quotient map $X\rightarrow X/A$ that collapses $A$ to a point is non-degenerate.
  3. (3)
    Let $P\rightarrow Q$ be a poset map. Then the induced map on order complexes [^9] Also known as the classifying space of the poset. $\Delta(X)\rightarrow\Delta(Y)$ is non-degenerate.
  4. (4)
    Let $I$ be the standard interval considered as a cell complex with two 0-cells that are connected by a 1-cell.
[…]
    Consider the map
    $f\colon I\rightarrow I,f(t)\coloneqq\begin{cases}t&t\leq\frac{1}{2}\\ 1-t&t>\frac{1}{2}\end{cases}.$
    Then $f$ is simplicial and even preserves the cover relation in $F(I)$, but $f$ is not non-degenerate.
    Similar degenerate collapse maps of a cell to one of its faces also exist in higher dimensions.
In the literature, there is an established notion of equivalence of CW complexes that takes both, combinatorial and topological features, into account:
Definition 3.21 ([FRA79, Definition 2.1][BH04, Definition 6.32]).
@@
The complexes $X$ and $X^{\prime}$ are called cell equivalent if and only if there is a homotopy equivalence $h\colon X\rightarrow X^{\prime}$ with the property that $h$ induces an isomorphism of face posets [^10] In [FRA79], the author originally mentioned a “one-to-on correspondence” of cells. We interpret the term “correspondence” in the more precise above-mentioned way of an induced map on face posets, which also seems to be what the author of [FRA79] intended. $F(h)\colon F(X)\xrightarrow{\cong}F(X^{\prime})$ such that if $e\subset X$ corresponds to $e^{\prime}\subset X^{\prime}$, then $h$ maps $X(e)$ to $X^{\prime}(e^{\prime})$ and restricts to a homotopy equivalence of these subcomplexes.
Remark 3.22.
It follows directly from the definition that cell equivalences are in particular simplicial and homotopy equivalences. For the context of discrete Morse theory, we prefer the point of view of simplicial maps between regular CW complexes because they focus more on the combinatorial structure of the face poset. However, in the context of arbitrary CW complexes one has to work with cell equivalences because they keep track of the irregularity of attaching maps.
Furthermore, cell equivalences induce isomorphisms of the corresponding face posets. This way, cell equivalences and non-degenerate maps are the link between geometric and abstract morphisms of regular CW complexes.
Lemma 3.23.
Let $\varphi\colon X\rightarrow Y$ be a simplicial map of regular CW complexes. If $\varphi$ preserves the cover relation $\prec$ in the corresponding face posets, then $\varphi$ maps $n$-cells to $n$-cells.
Proof.
We prove the statement by induction: Since simplicial maps are cellular, they map 0-cells to 0-cells. Let $\sigma$ be an $n$ cell of $X$. By inductive assumption, all n1n-1 cells $\sigma^{\prime}\subset\partial\sigma$ are mapped to n1n-1 cells of $Y$. Because $\sigma^{\prime}\prec\sigma$ and $\varphi$ preserves cover relations, we have $\varphi(\sigma^{\prime})\prec\varphi(\sigma)$. Because $Y$ is regular, this implies $\operatorname{dim}\varphi(\sigma)=n$. ∎
Proposition 3.24.
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map of regular CW complexes that preserves the cover relation. Then $\varphi$ induces homotopy equivalences on all subcomplexes $X(\sigma)$ for $\sigma$ any cell of $X$.
Moreover, if the induced map $F(\varphi)$ is injective, then $\varphi$ is a cell equivalence onto its image.
Proof.
We observe that for regular CW complexes $X$, the subcomplexes $X(\sigma)$ for an $n$-cell $\sigma$ are homeomorphic to $D^{n}$. Since $\varphi$ preserves cover relations, it also maps chains of cover relations to chains of cover relations. Since $\varphi$ is non-degenerative, for every cell $\sigma$ it preserves the initial vertices of all maximal chains of cover relations that terminate in $\sigma$. We prove inductively that for each such chain of cover relations $\sigma_{0}\prec\dots\prec\sigma_{i}\prec\dots\prec\sigma$ in $X$ that finishes in $\sigma\in X$, the restriction of $\varphi$ to each $\sigma_{i}$ is a cell equivalence onto its image.
For $i=0$, there is nothing to show. For $i=1$, it follows from preservation of cover relations and non-degeneracy, that any 1-cell $\sigma_{1}$ that is spanned by two 0-cells $\alpha,\beta$ needs to be mapped to a 1-cell that is spanned by $\varphi(\alpha)$ and $\varphi(\beta)$. Moreover, $\varphi$ is injective on the 0-cells of $\sigma$, because otherwise either
  1. (1)
    a 1-cell between two 0-cells would be mapped to a 0 cell, which would violate preservation of cover relations, or
  2. (2)
    such a 1-cell would be mapped to a degenerate 1-cell, i.e. both 0-cells of $\sigma_{1}$ would be mapped to the same 0-cell.
Thus, $\varphi_{\lvert X(\sigma)}$ is a continuous extension of a bijection $S^{0}\cong S^{0}$, and, hence, must be a homotopy equivalence.
Let $\sigma_{i}$ be an $i$-cell in the boundary of $\sigma$. By inductive assumption, $\varphi_{\lvert\partial\sigma_{i}}$ is a cell equivalence. In particular, $\varphi(\partial\sigma_{i})\simeq S^{n-1}$. Thus, preservation of cover relations and non-degeneracy imply that $\sigma_{i}$ is mapped to a cell $\tau\in Y$ such that the gluing map of $\tau$ is a homeomorphism onto $\varphi(\partial\sigma_{i})$, i.e. the map $\varphi_{\lvert X(\sigma_{i})}$ is a continuous extension of a cell equivalence $\partial\sigma_{i}\simeq\partial\tau$. This implies that $\varphi_{\lvert X(\sigma)}$ is a homotopy equivalence because $Y(\tau)\cong D^{n}$ is contractible.
For the second statement, we observe that if $F(\varphi)$ is injective, then we know that the induced map $F(\varphi)\colon F(X)\rightarrow F(\operatorname{im}(\varphi))$ is an isomorphism of face posets. Together with the first statement, it follows directly from Definition 3.21 that $\varphi\colon X\rightarrow\varphi(X)$ is a cell equivalence. ∎
Remark 3.25.
It turns out that all assumptions of Proposition 3.24 are necessary: If $\varphi$ is assumed to be non-degenerate but does not necessarily preserve cover relations, then $\varphi$ might be just a quotient map to a point.
If $\varphi$ is assumed to preserve cover relations but not necessarily to be non-degenerate, it might be a map as in Example 3.20 (4).
In both of these cases, $\varphi$ would not induce homotopy equivalences on subcomplexes of the form $X(\sigma)$.
Moreover, if we assume $\varphi$ to induce homotopies on subcomplexes of the form $X(\sigma$ but do not necessarily assume injectivity of the induced map on face posets, $\varphi$ could identify pairs of parallel cells as in the following example:
$\bullet\bullet\rightarrow\bullet\bullet$
In particular, the map $\varphi$ would not need to be a cell equivalence.
Having established the relationship between non-degenerate maps and cell equivalences, we want to use these notions to induce maps on spaces of discrete Morse functions. The straightforward way to do that is to use the contravariant functoriality of partially ordered coordinate spaces, see Remark 2.4.
Proposition 3.26.
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map of CW complexes. Then there is an induced linear map $\varphi^{*}\colon\mathbb{R}^{Y}\rightarrow\mathbb{R}^{X}$ given by $\varphi^{*}g\coloneqq g\circ\varphi$, for $g\colon Y\rightarrow\mathbb{R}$.
If $\varphi$ is injective, then $\varphi^{*{-1}}(\mathcal{A}(X))\subset\mathcal{A}(Y)$ is a subarrangement. Moreover, $\varphi^{*}$ restricts in the injective case to a map $\mathcal{M}(Y)\rightarrow\mathcal{M}(X)$.
Proof.
The first statement is true because non-degenerate maps induce maps on face posets. The second statement follows because non-degenerate maps preserve cover relations.
For the third statement we consider that injectivity implies that each pair of cells of $X$ that are related by a cover relation gets their function values from distinct cover relations in $Y$. Hence, the numbers from (1) and (2) of Definition 3.11 are preserved by $\varphi^{*}$. ∎
@@
We call the induced map $\varphi^{*}$ from Proposition 3.26 the pullback map induced by $\varphi$.
Since we require the relatively strong notion of non-degenerate maps, we also get a map in the other direction.
Lemma 3.28.
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map between CW complexes. Then $\varphi$ induces a linear map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ given by
$\varphi_{*}(f)(\sigma)\coloneqq\begin{cases}f(\varphi^{-1}(\sigma))&\text{ if }\lvert\varphi^{-1}(\sigma)\rvert=1\\ 0&else\end{cases}$
that maps $\mathcal{A}^{+}(X)$ to $\mathcal{A}^{+}(Y)$. If $\varphi$ is injective, then $\varphi_{*}$ embeds $\mathcal{A}^{+}(X)$ into $\mathcal{A}^{+}(Y)$.
@@
The first statement follows directly from the fact that $\varphi$ induces a map on face posets that preserves cover relations. The second statement follows because injectivity of $\varphi$ implies an injective correspondence of cover relations.
@@
In fact, if $\varphi$ is a degenerate simplicial map, $F(\varphi)$ does not preserve cover relations. Since $F(\varphi)$ is a poset map between finite posets, the fiber $F(\varphi)^{-1}(y)$ of a point $y\in F(Y)$ is either empty or an interval in $F(X)$. Intervals in $F(X)$ correspond to subspaces in $\mathbb{R}^{X}$ which means that $\varphi_{*}$ collapses the subspace corresponding to $F(\varphi)^{-1}(y)$ to one coordinate corresponding to $y$. Geometrically, this corresponds to collapsing111111in the sense of collapsing a topological space, not necessarily in the sense of simple collapses the corresponding subcomplex $U\subset X$ to a point.
@@
  • $\operatorname{Face}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\subset\sigma\text{ cover relation in }F(X)\}$, and
  • $\operatorname{Coface}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\supset\sigma\text{ cover relation in }F(X)\}$.